# HLSTester: Efficient Testing of Behavioral Discrepancies with LLMs for High-Level Synthesis

### v1

@@ Abstract.
In high-level synthesis (HLS), C/C++ programs with synthesis directives are used to generate circuits for FPGA implementations. However, hardware-specific and platform-dependent characteristics in these implementations can introduce behavioral discrepancies between the original C/C++ programs and the circuits after high-level synthesis. Existing methods for testing behavioral discrepancies in HLS are still immature, and the testing workflow requires significant human efforts. To address this challenge, we propose HLSTester, a large language model (LLM) aided testing framework that efficiently detects behavioral discrepancies in HLS. To mitigate hallucinations in LLMs and enhance prompt quality, the testbenches for original C/C++ programs are leveraged to guide LLMs in generating HLS-compatible testbenches, effectively eliminating certain traditional C/C++ constructs that are incompatible with HLS tools. Key variables are pinpointed through a backward slicing technique in both C/C++ and HLS programs to monitor their runtime spectra, enabling an in-depth analysis of the discrepancy symptoms. To reduce test time, a testing input generation mechanism is introduced to integrate dynamic mutation with insights from an LLM-based progressive reasoning chain. In addition, repetitive hardware testing is skipped by a redundancy-aware filtering technique for the generated test inputs. Experimental results demonstrate that the proposed LLM-aided testing framework significantly accelerates the testing workflow while achieving higher testbench simulation pass rates compared with the traditional method and the direct use of LLMs on the same HLS programs.
conference: Z; 2025; USA
1. Introduction
@@
As the demand for hardware performance in AI computing continues to grow, an increasing number of software applications originally designed for CPUs are being ported to specialized hardware accelerators like FPGAs (b1, ; b2, ; b3, ; b3.1, ). High-Level Synthesis (HLS) tools can take general-purpose programming languages such as C/C++ and generate Hardware Description Languages (HDL), thereby synthesizing corresponding circuits for FPGA deployment (b4.1, ; b4.2, ).
@@
Due to limitations of HLS tools, only a subset of C/C++ programs is supported, referred to as HLS programs (b5, ). When these programs are synthesized into circuits by the HLS tool, they may exhibit behavioral discrepancies compared to the original C/C++ programs running on CPUs. Such discrepancies include overflow or truncation from customized bit width, out-of-bounds memory accesses caused by static allocation, and data dependencies due to pipelining, etc. For example, default data types in CPU instructions are typically represented as int or float, whereas in FPGA designs, the bit width of integers and floats are often customized, which could lead to overflows on FPGAs. Additionally, FPGAs enable parallel computing through pipeline execution by incorporating hardware directives such as #pragma HLS pipeline, which may produce results that differ from those of sequential CPU execution, particularly in cases where data dependencies or feedback paths exist among functions. Therefore, testing behavioral discrepancies is important, as programs that execute correctly on CPUs may still fail or produce incorrect results when deployed on FPGAs (b7, ; b7.1, ; b7.3, ).
@@
To test behavioral discrepancies in HLS, developers typically need to monitor runtime spectra (e.g., the dynamic traces of variable values, accessed offsets, and stack/FIFO queue sizes) by constructing testbenches and instrumenting key variables. Any discrepancies in these spectra between C/C++ programs on CPUs and synthesized circuits on FPGAs can reveal issues in HLS that need to be addressed (b7.8, ; b7.9, ). However, this process requires developers to have expertise in both hardware and software, with significant efforts in the overall testing workflow. In addition, developers generally employ fuzz testing (b8, ), a common test input generation scheme that has been widely applied in software testing where each test input run completes in just a few milliseconds(b8.1, ). However, in hardware testing, particularly in FPGA testing where clock-cycle-level simulation is required due to data dependencies, the testing latency (i.e., the time required to perform a test operation) often ranges from several minutes to several hours, rather than milliseconds(b8.2, ; b8.3, ). Moreover, test inputs generated by fuzz testing do not take hardware platform characteristics into account, making it ineffective in testing behavioral discrepancies in HLS.
@@
Large language models (LLMs) have shown great potential in automated software and hardware design, thereby supporting the entire process from initial concepts to design, testing, verification, and physical implementation (b10, ; b11, ; b11.1, ; b11.2, ; b11.3, ; b11.7, ; b11.8, ; b11.9, ). Recent research has shown that LLMs can be employed to test errors in C/C++ programs and Hardware Description Language (HDL) designs. CityWalk (b12, ) leverages project-dependency analysis and insights from documentation to address complex features, such as pointers, templates, and virtual functions, thus producing high-quality testbenches. CorrectBench (b13, ) employs LLMs to implement a testing platform with self-correction, and an evaluation framework is proposed to evaluate the performance of the generated testbenches. VerilogReader (b14, ) leverages LLMs to interpret Verilog code, aiming to generate code convergence tests. A testing interpreter is introduced to enrich prompts and enhance LLM’s understanding of the testing intent.
@@
Previous methods above use LLMs to test simple errors in common code types, such as individual C/C++ or Verilog programs that LLMs are extensively trained on and thus highly familiar with. However, these methods do not provide a comprehensive testing workflow to detect discrepancy symptoms in interdisciplinary fields such as HLS, which require both hardware and software exploration. In this paper, an LLM-aided HLS testing framework, HLSTester, is proposed that employs an iterative testing workflow to automatically detect behavioral discrepancies in the HLS design. The key contributions of this paper are summarized as follows.
@@
Table 1.  Typical sources of discrepancies when porting C/C++ programs from CPU to FPGA via High-Level Synthesis.
@@
  • We propose an efficient testing framework for HLS that covers code analysis, testbench generation, and runtime profiling for behavioral discrepancy detection.
@@
  • By leveraging the adaptive learning capability of LLMs, existing C/C++ testbenches are used as references to guide the LLM to generate testbenches that are compatible with the HLS tool, which improves the simulation pass rate of the generated testbenches by an average of 20.67%.
@@
  • To enable an in-depth analysis of the discrepancy, a backward slicing technique, combined with the semantic understanding capabilities of LLMs, is employed to instrument the HLS program. Monitoring scripts are also developed to record the runtime spectra of key variables, which could also provide a reliable basis for subsequent test input mutation.
@@
  • To further reduce the latency of HLS testing, a redundancy-aware technique is proposed to identify and skip repetitive hardware simulations among the generated test inputs, which contributes an additional 15.73% acceleration in the testing workflow.
@@
The rest of this paper is organized as follows. Section 2 provides the background and motivation for this work. Section 3 details the proposed framework. The experimental results are provided in Section 4. Section 5 concludes the paper.
2. Background and Motivation
@@
Since some traditional C/C++ constructs are not supported by HLS tools or lead to inefficient synthesized circuits, significant code refactoring and the incorporation of specific hardware directives are necessary to adapt them for synthesis and optimization (b5, ). Testing behavioral discrepancies between the original C/C++ programs and the circuits after high-level synthesis is challenging due to assumptions made during hardware synthesis. Table 1 presents typical examples of how directives cause these discrepancies (b6, ).
@@
$\circ$ Data Type: HLS supports arbitrary bit width for variables, leading to a reduction in FPGA resource usage. For example, analysis of the variable ‘m’ in a standard C++ program exhibits a minimum value of 0 and a maximum of 481, thereby necessitating only 9 bits instead of the 32-bit int type. In such cases, the directive #typedef ap_uint$<$9$>$ is used to declare a 9-bit arbitrary precision integer. With custom bit width, FPGAs become more susceptible to overflow, potentially leading to incorrect results.
Figure 1.  Traditional workflow for testing behavioral discrepancies in High-Level Synthesis.
@@
$\circ$ Memory Management: HLS tools do not support dynamic arrays because FPGAs are unable to manage unbounded data structures. As a result, functions such as malloc() should be replaced with pre-allocated static arrays of estimated size (b14.3, ). Accessing an unexpected address may cause the FPGA to return an incorrect output.
@@
$\circ$ Parallelization and Loop: Hardware directives such as #pragma HLS pipeline can be inserted within loops to enable pipelining, allowing multiple input streams to be processed concurrently within the same clock cycle, thereby improving data throughput. The directive #pragma HLS unroll creates multiple instances of a loop body in the generated RTL design, enabling partial or complete parallel execution of loop iterations. In addition, the directive #pragma HLS dataflow enables task-level pipelining, allowing functions and loops to overlap their operations (b14.4, ). However, such parallel execution can produce results that differ from sequential CPU execution, particularly when misaligned data accesses occur or feedback paths exist between different variables or functions.
@@
$\circ$ Recursion: HLS tools do not support dynamic stack allocation for execution states, so recursion needs to be converted to iteration using a finite-size stack (b14.5, ). When recursion depth exceeds the pre-estimated size, FPGA compilation failures occur.
@@
An overview of the traditional workflow for testing behavioral discrepancies in HLS is shown in Fig. 1. Typically, developers begin with a program analysis, followed by modifications to the testbench to eliminate certain traditional C/C++ constructs that are incompatible with HLS tools. Then, both the original C/C++ programs and their HLS counterparts are instrumented, and a set of relevant test inputs is designed. At runtime, key spectra values (i.e., variable traces) are recorded and compared between the original C/C++ program and the circuits after high-level synthesis, and any mismatches in these spectra indicate behavioral discrepancies.
To generate these test inputs in the HLS workflow, several methods are typically employed: (1) random generation using data generators (b15, ); (2) handcrafted inputs by an HLS expert; or (3) systematic enumeration through scripts (b16, ). Typically, the test input generation technique begins with a set of seed inputs, selecting one seed randomly and producing new inputs by altering a few bits or bytes. However, this method relies on the assumption that the target program completes execution in milliseconds during software testing. However, in FPGA testing, a single hardware synthesis and testing simulation may take several minutes or even hours. Moreover, traditional mutation techniques in software testing overlook hardware-specific characteristics such as custom bit widths, static memory allocation, and pipelining via hardware directives, resulting in ineffective behavioral discrepancy testing in HLS.
Figure 2.  Example of behavioral discrepancies between (a) the original C++ code and (b) the HLS code.
@@
To illustrate the need for testing behavioral discrepancies in HLS environments, Fig. 2(a) presents an edge processing task. In this task, the original C++ program first computes the cumulative sum of an image gradient vector and then compares the percentage of each gradient against a threshold to decide whether to process pixels. In Fig. 2 (b), to address the performance bottleneck on the CPU and fully exploit FPGA parallelism, the developer refactored the C++ code, which introduced custom bit width, static arrays, and various hardware directives. However, these hardware-specific refactoring may induce platform-dependent behavioral discrepancies. For example, custom bit width could result in overflow, division-by-zero, or truncation issues; replacing dynamic memory allocation with static arrays may lead to out-of-bound accesses; and the pipelining may produce results that differ from sequential CPU execution.
@@
Therefore, it is necessary to develop an efficient testing framework tailored for HLS. Contrary to previous work, the proposed HLSTester leverages the in-context learning and comprehension capabilities of LLMs to assist key testing points. Integrating testing scripts with feedback from HLS tools accelerates discrepancy testing and results in a more efficient testing workflow.
3. LLM-Aided Behavior Discrepancy Testing for High-Level Synthesis
@@
To efficiently detect behavioral discrepancies between the original C/C++ programs and the circuits after high-level synthesis, we propose an LLM-aided testing framework composed of five stages: preprocessing, code instrumentation, spectra monitoring, test input generation, and redundant test input filtering. As illustrated in Fig. 3, the framework first constructs a testbench compatible with the HLS tool by eliminating incompatible C/C++ constructs. Next, a backward slicing technique is employed to pinpoint key variables, which are then instrumented to enable real-time monitoring of their runtime spectra. The collected spectra are fed back into the test input generation stage to guide the dynamic mutation of the test inputs, forming a cyclic feedback system. Finally, a redundancy-aware filtering mechanism for the generated test inputs is applied to skip duplicate hardware simulations. Each stage will be illustrated in detail in the following subsections.
Figure 3. Testing Workflow of HLSTester.
3.1. Preprocessing
Certain C/C++ constructs in C/C++ testbenches, such as traditional C/C++ header files, definitions for dynamic arrays, and other unsupported features, are incompatible with HLS tools. However, despite the strict constraints imposed by HLS tools, both types of testbenches exhibit notable similarities in their intended functionality and basic structure. To enable the generation of a testbench compatible with HLS tools, three key components are prepared: (1) the original C/C++ program, (2) its corresponding C/C++ testbench, and (3) a refactored version of the program, referred to as the HLS program under test, which complies with the constraints of the HLS tool. At this stage, given the adaptive learning capabilities of LLMs, existing C/C++ testbenches are used as references to guide the generation of testbenches suited for HLS tools. We formalize this HLS-compatible testbench generation as follows:
(1) $F_{\mathrm{LLM}}:T_{\mathrm{C/C++}}\;\;\to\;\;T_{\mathrm{HLS}},\quad\text{s.t.}\quad\mathrm{Constraints}\_{\mathrm{HLS},}\vspace{-0.17cm}$
where $T_{\mathrm{C/C++}}$ is the set of traditional C/C++ testbenches, $T_{\mathrm{HLS}}$ is the set of HLS-compatible testbenches, $\mathrm{Constraints}\_{\mathrm{HLS}}$ are the constraints imposed by HLS tools. The LLM needs to focus on the following aspects during the analysis and generation process:

### v3

@@ Abstract
In high-level synthesis (HLS), C/C++ programs with synthesis directives are used to generate circuits for FPGA implementations. However, hardware-specific and platform-dependent characteristics in these implementations can introduce behavioral discrepancies between the original C/C++ programs and the circuits after high-level synthesis. Existing methods for testing behavioral discrepancies in HLS are still immature, and the testing workflow requires significant human efforts. To address this challenge, we propose HLSTester, a large language model (LLM) aided testing framework that efficiently detects behavioral discrepancies in HLS. To mitigate hallucinations in LLMs and enhance prompt quality, existing C/C++ testbenches are used to guide the LLM to generate HLS-compatible versions, effectively eliminating certain traditional C/C++ constructs that are incompatible with HLS tools. Key variables are pinpointed through a backward slicing technique in both C/C++ and HLS programs to monitor their runtime spectra, enabling an in-depth analysis of the discrepancy symptoms. Then, a testing input generation mechanism is introduced to integrate dynamic mutation with insights from an LLM-based progressive reasoning chain. In addition, repetitive hardware testing is skipped by a redundancy-aware technique for the generated test inputs. Experimental results demonstrate that the proposed LLM-aided testing framework significantly accelerates the testing workflow while achieving higher testbench simulation pass rates compared with the traditional method and the direct use of LLMs on the same HLS programs.
I Introduction
@@
As the demand for hardware performance in AI computing continues to grow, an increasing number of software applications originally designed for CPUs are being ported to specialized hardware accelerators like FPGAs [1, 2, 3]. High-Level Synthesis (HLS) tools can take general-purpose programming languages such as C/C++ and generate Hardware Description Language (HDL) designs, thereby synthesizing corresponding circuits for FPGA deployment [4, 5, 6].
@@
Due to limitations of HLS tools, only a subset of C/C++ programs is supported, referred to as HLS programs [7]. When these programs are synthesized into circuits by the HLS tool, they may exhibit behavioral discrepancies compared with the original C/C++ programs running on CPUs. Such discrepancies include overflow from customized bit width, out-of-bounds memory accesses caused by static allocation, data dependencies due to pipelining, etc. For example, default data types on CPUs are typically represented as int or float, whereas in FPGA designs, the bit width of integers and floats are often customized, which could lead to overflows on FPGAs. Additionally, FPGAs enable parallel computing through pipeline execution by incorporating hardware directives such as #pragma HLS pipeline, which may produce results that differ from those of sequential CPU execution, particularly in cases where data dependencies or feedback paths exist among functions.
@@
Given these potential divergences, systematic testing is essential in the HLS workflow, as programs that execute correctly on CPUs may still fail or produce incorrect results when deployed on FPGAs [9]. To test behavioral discrepancies in HLS, developers typically need to monitor runtime spectra by constructing testbenches and instrumenting key variables. The term spectra refers to dynamic traces that capture variable values, memory access patterns, stack sizes, FIFO queue sizes, and other relevant runtime metrics. Any discrepancies in these spectra between C/C++ programs on CPUs and synthesized circuits on FPGAs can reveal issues in HLS that need to be addressed [10]. However, this testing process requires developers to have expertise in both hardware and software, with significant efforts in the overall testing workflow. In addition, developers generally employ fuzz testing [11], a common test input generation scheme that has been widely applied in software testing where each test run completes in just a few milliseconds[12]. However, in hardware testing, particularly in FPGA testing where clock-cycle-level simulation is required due to data dependencies, the testing latency (i.e., the time required to perform a test operation) often ranges from several minutes to several hours, rather than milliseconds[13]. Moreover, test inputs generated by fuzz testing do not take hardware platform characteristics into account, making it ineffective in testing behavioral discrepancies in HLS.
@@
To overcome these limitations, more advanced testing methodologies are emerging. Large language models (LLMs) have demonstrated great potential in automated software and hardware tasks, thereby supporting the entire process from initial design to testing, verification, and physical implementation [14, 15, 16, 17, 18, 19, 20]. Recent research has shown that LLMs can be employed to test errors in C/C++ programs and Hardware Description Language (HDL) designs. CityWalk [23] leverages project-dependency analysis and insights from documentation to address complex features, such as templates and virtual functions, thus producing high-quality testbenches. CorrectBench [24] employs LLMs to implement a testing platform with self-correction, and an automatic validation framework is proposed to evaluate the performance of the generated testbenches.
@@
Previous methods above use LLMs to test simple errors in common code types, such as individual C/C++ or Verilog programs that LLMs are extensively trained on and thus highly familiar with. However, these methods do not provide a testing workflow to detect behavioral discrepancies in interdisciplinary fields such as HLS, which necessitates the co‑exploration of hardware and software [14]. In this paper, we introduce HLSTester, an LLM‑aided HLS testing framework that employs an iterative testing workflow to automatically detect behavioral discrepancies in the HLS design. The key contributions of this paper are summarized as follows.
@@
TABLE I: TYPICAL DISCREPANCY SOURCES WHEN PORTING C/C++ PROGRAMS FROM CPU TO FPGA VIA HIGH-LEVEL SYNTHESIS.
@@
  • We propose an efficient testing framework for HLS that covers code analysis, testbench modification, and runtime profiling for behavioral discrepancy detection.
@@
  • By leveraging the adaptive learning capability of LLMs, existing C/C++ testbenches are used to guide the LLM to generate HLS-compatible versions, improving the pass rate of the resulting testbenches by an average of 20.67%.
@@
  • To enable an in-depth analysis of the discrepancy, a backward slicing technique, combined with the semantic understanding capabilities of LLMs, is employed to instrument the HLS program. Monitoring scripts are also developed to record the runtime spectra of key variables, which could also provide a basis for test input mutation.
@@
  • To avoid redundant hardware simulations during HLS testing, a redundancy-aware technique is proposed to identify and skip repetitive test inputs, which contributes an additional 15.73% acceleration in the testing workflow.
@@
The rest of this paper is organized as follows. Section II provides the background and motivation for this work. Section III details the proposed framework. The experimental results are provided in Section IV. Section V concludes the paper.
II Background and Motivation
@@
Since some traditional C/C++ constructs are not supported by HLS tools or lead to inefficient synthesized circuits, significant code refactoring and the incorporation of specific hardware directives are necessary to adapt them for synthesis[7]. Testing behavioral discrepancies between the original C/C++ programs and the circuits after high-level synthesis is challenging due to assumptions made during synthesis. Table I presents typical examples of how directives cause these discrepancies [8].
@@
$\circ$ Data Type: HLS supports arbitrary bit width for variables, leading to a reduction in FPGA resource usage. For example, analysis of the variable ‘m’ in a standard C++ program exhibits a minimum value of 0 and a maximum of 481, thereby necessitating only 9 bits instead of the 32-bit int type. In such cases, the directive #typedef ap_uint$<$9$>$ is used to declare a 9-bit precision integer. With custom bit width, FPGAs become more susceptible to overflow, leading to incorrect results.
Figure 1:  Traditional workflow for testing behavioral discrepancies in HLS.
@@
$\circ$ Memory Management: HLS tools do not support dynamic arrays because FPGAs are unable to manage unbounded data structures. As a result, functions such as malloc() should be replaced with static arrays. Accessing an unexpected address may cause the FPGA to return an incorrect output.
@@
$\circ$ Loop and Parallelization: Hardware directives such as #pragma HLS unroll create multiple instances of a loop body in the generated RTL design, enabling parallel execution of loop iterations. In addition, the directive #pragma HLS dataflow enables task-level pipelining, allowing functions to overlap their operations [25]. However, such parallel execution can produce results that differ from sequential CPU execution, particularly when misaligned data accesses occur, or feedback paths exist between different variables or functions.
@@
$\circ$ Recursion: HLS tools do not support dynamic stack allocation for execution states, so recursion needs to be converted to iteration using a finite-size stack [26]. When recursion depth exceeds the estimated size, FPGA compilation failures occur.
@@
An overview of the traditional workflow for testing behavioral discrepancies in HLS is shown in Fig. 1. Typically, developers begin with a program analysis, followed by modifications to the testbench to eliminate certain traditional C/C++ constructs that are incompatible with HLS tools. Then, both the original C/C++ programs and their HLS counterparts are instrumented, and a set of relevant test inputs is designed to be applied identically in both software and hardware simulations. At runtime, spectra of key variables are recorded and compared between the original C/C++ program and the circuits after high-level synthesis, and any mismatches in these spectra indicate behavioral discrepancies. On the right side of Fig. 1, such discrepancies may arise from overflows, truncations, and issues introduced by other hardware directives, which are then used to debug the refactored HLS program.
Figure 2:  Discrepancies between (a) the original C++ code & (b) the HLS code.
@@
To illustrate the need for testing behavioral discrepancies in HLS environments, Fig. 2(a) presents an edge processing task. In this task, the original C++ program first computes the cumulative sum of an image gradient vector and then compares the percentage of each gradient against a threshold to decide whether to process pixels. In Fig. 2 (b), to address the performance bottleneck on the CPU and fully exploit FPGA parallelism, the developer refactored the C++ code, which introduced custom bit width, static arrays, and various hardware directives. However, these hardware-specific refactorings may induce platform-dependent behavioral discrepancies. For example, custom bit width could result in overflow or truncation issues; replacing dynamic memory allocation with static arrays may lead to out-of-bound accesses; and the pipelining may produce results that differ from sequential CPU execution.
To generate these test inputs in the HLS workflow, several methods are typically employed: (1) random generation using data generators [27]; (2) handcrafted inputs by an HLS expert; or (3) enumeration through scripts [28, 29, 30]. Typically, the test input generation begins with a set of seed inputs, selecting one seed randomly and producing new inputs by altering a few bits. However, in FPGA testing, a single hardware synthesis and testing simulation may take several minutes or even hours. Moreover, traditional mutation techniques in software testing overlook hardware-specific characteristics such as custom bit width, static allocation, and pipelining via hardware directives, resulting in ineffective behavioral discrepancy testing in HLS.
@@
Contrary to previous work, this paper introduces HLSTester, a framework that leverages the in-context learning and comprehension capabilities of LLMs to assist key testing points. Integrating testing scripts with feedback from HLS tools accelerates discrepancy testing and leads an efficient testing workflow.
III LLM-Aided Behavioral Discrepancy Testing
@@
The proposed testing framework is composed of five stages: (1) testbench modification for HLS, (2) code instrumentation, (3) runtime spectra monitoring, (4) test input generation, and (5) redundancy-aware filtering. As illustrated in Fig. 3, the framework first constructs a testbench compatible with the HLS tool by eliminating incompatible C/C++ constructs. Next, a backward slicing technique is employed to pinpoint key variables, which are then instrumented to enable monitoring of their runtime spectra. The collected spectra are fed back into the test input generation stage to guide the dynamic mutation of the test inputs, forming a cyclic feedback system. Finally, a redundancy-aware technique for the generated test inputs is applied to skip duplicate hardware simulations. Each stage will be described in detail in the following subsections. In the following description, we use the terms below in the HLS flow:
  • Original C/C++ program: The original software implementation written in C/C++ designed for CPU execution.
  • HLS program: The HLS-compatible program converted from the original C/C++ program with directives.
  • Original testbench: The original C/C++ testbench developed to test the original C/C++ program.
  • HLS-compatible testbench: A testbench modified from the original C/C++ testbench and able to be compiled by HLS tools to test the generated RTL design.
Figure 3: Testing Workflow of the proposed HLSTester.
III-A Testbench Modification for HLS
To test the discrepancies between a given C/C++ program and the HLS program converted from it, a testbench is needed. However, the original C/C++ testbench cannot be used directly, as it may contain C/C++ constructs that are incompatible with HLS tools and thus cannot be compiled to test the target design, e.g., traditional C/C++ header files, definitions for dynamic arrays, and other unsupported features. To deal with this problem, we use an LLM to modify the original C/C++ testbench to generate the HLS-compatible testbench. The LLM focuses on the following aspects during the analysis and modification process to generate the HLS-compatible testbench: